OpenStudy (anonymous):

A laser produces 20.0 mW of red light. In 1.00 hr, the laser emits 2.57×1020 photons. What is the wavelength of the laser?

OpenStudy (ookawaiioo):

0.020 Watts * 3600sec = Energy Energy = h*f = h*(c/wavelength) Wavelength = h*c/Energy h = planck constant (I forgot what the values were) c = speed of light (3.0 x 10^8 m/s) Goodluck